OpenStudy (anonymous):

Write the equation of the line with a slope of 3/7 passing through the point (9, -2).

OpenStudy (katrinakaif):

Use the equation: (y - y1) = m (x - x1) m = slope = 3/7 (x1,y1) = (9, -2) And solve =)

OpenStudy (anonymous):

im confused..

OpenStudy (katrinakaif):

Okay lets walk thru it step by step: Follow below =)

OpenStudy (katrinakaif):

To find the equation using the slope and an ordered pair we use the equation: (y - y1) = m (x - x1)

OpenStudy (anonymous):

ok..

OpenStudy (katrinakaif):

Now the next step is the PLUG in what you need to plug in. Since you already have the slope, which is "m" in the equation, plug it into the equation: (y-y1) = 3/7 (x-x1)

OpenStudy (anonymous):

alright..

OpenStudy (katrinakaif):

Next plug in the ordered pairs..Do you know where they go?

OpenStudy (anonymous):

x and y?

OpenStudy (katrinakaif):

Not x and y but as (x1,y1) = ( 9,-2)

OpenStudy (anonymous):

ok...

OpenStudy (katrinakaif):

So lets plug it all in as we have so far: (y - (-2)) = 3/7 (x - 9) The negative two (-2) becomes positive because its already negative and we are also substracting so + 2 ==> y + 2 = 3/7 (x - 9) So far?

OpenStudy (anonymous):

ok...

OpenStudy (katrinakaif):

Now simply convert this into a y = mx + b format Distribute 3/7 to (x - 9) = y + 2 = 3/7x - 27/7

OpenStudy (anonymous):

soo y=3/7x+9?

OpenStudy (katrinakaif):

Where did you get + 9 from? O_o

OpenStudy (anonymous):

(9,-2) ?? im really not sure

OpenStudy (katrinakaif):

Well okay: We plugged in the values to get: (y + y1) = 3/7 ( x -x1) (y +2) = 3/7 ( x - 9) Now distribute 3/7 to "x" and "-9" getting: (y + 2) = 3/7x - 27/7 How did i get 27/7? Well I had to distribute correct? and mulitply - 9 with 3/7 ? that means -9 x 3 _______ 7 which is equal to -27/7 So far we have: y + 2 = 3/7x - 27/7

OpenStudy (anonymous):

so y=3/7x-2?

OpenStudy (katrinakaif):

Nope..I ll explain

OpenStudy (anonymous):

?

OpenStudy (katrinakaif):

y + 2 = 3/7x - 27/7 Bring the 2 to the other side by substracting it from -27/7 Solve this for me: -27 2 _____ - ___ 7 1

OpenStudy (katrinakaif):

When I did that I got -25/7 SO final equation would be: y = 3/7x - 25/7 =)

OpenStudy (anonymous):

well it says the answer is either: a. 2x-9y=-3/7 b.y=3/7x+9 c.y=3/7x-2 d.3x-7y=41

OpenStudy (katrinakaif):

Hmm its either b or c... thinking

OpenStudy (anonymous):

ok

hero (hero):

What the...

OpenStudy (katrinakaif):

Oh I beleive its c because the y intercept ("b") is supposed to be -2 in the equation y = mx + b

OpenStudy (anonymous):

ok.. thanx!

OpenStudy (katrinakaif):

nO pROBLEM1

hero (hero):

What? I don't think you did that right. You were supposed to subtract 2, not assume it's the y intercept

OpenStudy (katrinakaif):

-2 is given as Y1 in the ordered paiir (x1,y1) (9, -2)

hero (hero):

You subtracted 2 incorrectly, the y-intercept is not 25/7

OpenStudy (anonymous):

its not c?

hero (hero):

I'm pretty sure the answer is d

hero (hero):

for not so obvious reasons. Looks like someone forgot how to subtract fractions

hero (hero):

Before you subtract a whole number from a fraction, you have to remember that the whole number must be converted to a fraction first.

OpenStudy (katrinakaif):

Forget hte math Hero because she doesn't need the answer in y = mx + b format..all you have to do is plug in for x and y? sigh idk

OpenStudy (anonymous):

is it still c??

OpenStudy (katrinakaif):

Sigh I am not sure..I believe it C but I am not confident..Hero?

hero (hero):

The answer is d

hero (hero):

If you pick c, you will be wrong

OpenStudy (anonymous):

im REALLY confused.. are you sure its d?

hero (hero):

If you're confused its because you don't know how to properly subtract fractions